# HUD: Hierarchical Uncertainty-Aware Disambiguation Network for Composed Video Retrieval

### v1

@@ Appendix A Details of Datasets
For a comprehensive evaluation, we select four datasets, which include one CVR dataset, WebVid-CoVR (Ventura et al., 2024), and two CIR datasets, FashionIQ (Wu et al., 2021) and CIRR (Liu et al., 2021). We now describe each dataset in detail as follows.
@@ A.1. CVR datasets
WebVid-CoVR (Ventura et al., 2024) serves as the first large-scale benchmark tailored for the CVR task. This dataset is generated by processing the WebVid-2M(Bain et al., 2021) dataset, yielding million CVR triplets that span unique videos and distinct modification texts. On average, each video lasts around seconds, and each modification text consists of approximately words. Moreover, each target video is associated with roughly triplets. The test set of WebVid-CoVR includes high-quality triplets, which were carefully chosen from the separate WebVid-10M (Bain et al., 2021) dataset after an intensive process of annotation and noise removal, resulting in a robust and challenging evaluation benchmark.
@@ A.2. CIR datasets
FashionIQ(Wu et al., 2021) is a widely used retrieval dataset in the fashion domain for CIR tasks, and the images are sourced from Amazon.com. The dataset comprises images split into three categories: Dresses, Shirts, and Tops&Tees. Following the evaluation protocol outlined in (Wu et al., 2021), we treat each category as a separate dataset for our experiments. Consistent with previous CIR approaches, approximately images are allocated for training and around for testing, resulting in about triplets for training and roughly triplets for testing.
CIRR(Liu et al., 2021) is an open-domain dataset introduced for the CIR task. It consists of real images sourced from the well-known language reasoning dataset (Suhr et al., 2019), which has been widely applied in natural language reasoning studies. The CIRR dataset comprises a total of triplets, with allocated for training, for validation, and the remaining for testing. Additionally, CIRR features a specialized subset designed for fine discrimination. This subset includes negative images that are visually very similar, thereby facilitating the evaluation of the model’s ability to distinguish between subtle false negatives.
@@ Appendix C Probabilistic Embedding Visualization
In order to intuitively analyze the role of the probabilistic embedding modeled by HUD, we employ t-SNE (Van der Maaten and Hinton, 2008) to visualize the relative positions of various features in the embedding space, as shown in Figure 6. In the figure, different colored points represent the visual detail embedding, textual probabilistic embedding, original composition, probabilistic composition(with probabilistic embedding), as well as the reference, target, and modification tokens. The green and blue ellipses represent the variance ranges for the textual and visual modalities, respectively. From the visualization, we obtain the following observations.
@@ Appendix E More Related Work
Large Language Models for Multimodal Retrieval. The advent of Large Language Models (Xin et al., 2025a; Wang et al., 2024; Zhou et al., 2024; Chu et al., 2025c; Zhang et al., 2025a; Zeng et al., 2025; Xin et al., 2025b) (LLMs) has significantly revolutionized the field of vision-language understanding (Qian et al., 2025a; Liu et al., 2024b; Zhang et al., 2025b; Liu et al., 2025d; Qian et al., 2025b; Lu et al., 2025a; Zhou et al., 2025). In the context of retrieval tasks, LLMs contribute primarily by enhancing the comprehension of the query text. For Composed Video Retrieval (CVR), the modification text often involves complex logic (e.g., “change object A to B but keep the background C”). LLM-based architectures (Shen et al., 2025; Chu et al., 2025a, b; Zhang et al., 2024b; Li et al., 2024; Xin et al., 2024c) can leverage their immense pre-trained knowledge to perform reasoning on these instructions (Luo et al., 2025; Sha et al., 2024; Kong et al., 2024, 2025), thereby extracting richer query representations (Cui et al., 2025; Sha et al., 2022; Li et al., 2025; Yi et al., 2024).
However, deploying full-scale LLMs for the retrieval of fine-grained visual details remains challenging, particularly in distributed or privacy-sensitive scenarios such as Federated Learning (FL) (Liu et al., 2025b). First, the sheer model size and inference latency of LLMs are often prohibitive for FL settings (Liu et al., 2025a), where models must be lightweight enough to run on resource-constrained edge devices while minimizing communication overhead (Zhao et al., 2025; Xin et al., 2024a, b, d; Lu et al., 2025c). Second, while LLMs excel at general semantic reasoning (Xiang et al., 2025; Liu et al., 2025c; Xu et al., 2025; Zhang et al., 2024a; Lu et al., 2025b; Liu et al., 2024a), they may struggle to adapt to the statistical heterogeneity of local client data or lack the granularity required to ground “atomistic” visual details (such as the small “chalk” or “window” instances discussed in our case studies) without centralized aggregation. Our HUD framework addresses these limitations through a lightweight hierarchical design. Instead of relying on the heavy computation of end-to-end Video-LLMs, HUD explicitly models uncertainty and incorporates an efficient atomistic-level alignment module. This ensures that the model captures subtle visual modifications and is amenable to deployment in decentralized environments, balancing precise retrieval with the efficiency required for potential federated applications.

### v2

@@ Appendix A Details of Datasets
For a comprehensive evaluation, we select four datasets, which include one CVR dataset, WebVid-CoVR (covr), and two CIR datasets, FashionIQ (FashionIQ) and CIRR (cirr). We now describe each dataset in detail as follows.
@@ A.1. CVR datasets
WebVid-CoVR (covr) serves as the first large-scale benchmark tailored for the CVR task. This dataset is generated by processing the WebVid-2M(Webvid-2M) dataset, yielding million CVR triplets that span unique videos and distinct modification texts. On average, each video lasts around seconds, and each modification text consists of approximately words. Moreover, each target video is associated with roughly triplets. The test set of WebVid-CoVR includes high-quality triplets, which were carefully chosen from the separate WebVid-10M (Webvid-2M) dataset after an intensive process of annotation and noise removal, resulting in a robust and challenging evaluation benchmark.
@@ A.2. CIR datasets
FashionIQ(FashionIQ) is a widely used retrieval dataset in the fashion domain for CIR tasks, and the images are sourced from Amazon.com. The dataset comprises images split into three categories: Dresses, Shirts, and Tops&Tees. Following the evaluation protocol outlined in (FashionIQ), we treat each category as a separate dataset for our experiments. Consistent with previous CIR approaches, approximately images are allocated for training and around for testing, resulting in about triplets for training and roughly triplets for testing.
CIRR(cirr) is an open-domain dataset introduced for the CIR task. It consists of real images sourced from the well-known language reasoning dataset (NLVR2), which has been widely applied in natural language reasoning studies. The CIRR dataset comprises a total of triplets, with allocated for training, for validation, and the remaining for testing. Additionally, CIRR features a specialized subset designed for fine discrimination. This subset includes negative images that are visually very similar, thereby facilitating the evaluation of the model’s ability to distinguish between subtle false negatives.
@@ Appendix C Probabilistic Embedding Visualization
In order to intuitively analyze the role of the probabilistic embedding modeled by HUD, we employ t-SNE (tsne) to visualize the relative positions of various features in the embedding space, as shown in Figure 6. In the figure, different colored points represent the visual detail embedding, textual probabilistic embedding, original composition, probabilistic composition(with probabilistic embedding), as well as the reference, target, and modification tokens. The green and blue ellipses represent the variance ranges for the textual and visual modalities, respectively. From the visualization, we obtain the following observations.
@@ Appendix E More Related Work
Large Language Models for Multimodal Retrieval. The advent of Large Language Models (xin2025lumina; wang2024enhancing; zhou2024human; chu2025dynamic; zhang2025rearank; zeng2025bridging; xin2025luminamgpt) (LLMs) has significantly revolutionized the field of vision-language understanding (qian2025adaptive; liu2024rag; zhang2025assessing; liu2025qfft; qian2025dyncim; lu2025journey; zhou2025glimpse). In the context of retrieval tasks, LLMs contribute primarily by enhancing the comprehension of the query text. For Composed Video Retrieval (CVR), the modification text often involves complex logic (e.g., “change object A to B but keep the background C”). LLM-based architectures (shen2025expertflow; chu2025selective; chu2025mcam; zhang2024seppo; li2024synergized; xin2024v) can leverage their immense pre-trained knowledge to perform reasoning on these instructions (luo2025tr; sha2024hierarchical; kong2024tiger; kong2025accelerating), thereby extracting richer query representations (11197022; sha2022regional; li2025hyfacialhybridfeatureextraction; yi2024towards).
However, deploying full-scale LLMs for the retrieval of fine-grained visual details remains challenging, particularly in distributed or privacy-sensitive scenarios such as Federated Learning (FL) (liu2025consistency). First, the sheer model size and inference latency of LLMs are often prohibitive for FL settings (liuimproving), where models must be lightweight enough to run on resource-constrained edge devices while minimizing communication overhead (zhao2025recondreamer++; xin2024vmt; xin2024mmap; xin2024parameter; lu2025differentiable). Second, while LLMs excel at general semantic reasoning (xiang2025promptsculptor; liu2025fedadamw; xu2025mmm; zhang2024building; lu2025causalsr; liu2024fedbcgd), they may struggle to adapt to the statistical heterogeneity of local client data or lack the granularity required to ground “atomistic” visual details (such as the small “chalk” or “window” instances discussed in our case studies) without centralized aggregation. Our HUD framework addresses these limitations through a lightweight hierarchical design. Instead of relying on the heavy computation of end-to-end Video-LLMs, HUD explicitly models uncertainty and incorporates an efficient atomistic-level alignment module. This ensures that the model captures subtle visual modifications and is amenable to deployment in decentralized environments, balancing precise retrieval with the efficiency required for potential federated applications.